• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT
    Make New Entry, Make Followup Entry

    User name Markowitz

    Log entry time 08:00:58 on June 26, 2010

    Entry number 328344

    keyword=Owl shift summary

    00:00 MCC is asked to deliver beam; MCC starts their procedure (BCM calibrations, check beam on dump, BPM check, etc.). MCC expects it to be "a little while" before the hall can start the plan.

    00:30 Looking at cosmic runs.

    01:00 MCC update. Magnet power supplies in the BSY need to be replaced and the MCC is now doing that. The beam "will be a while".

    02:35 MCC update. MCC has beam to dump and is now preparing to execute the MCC commissioning of the beam. It will be "at least another hour" before the beam is turned over to the experiment.

    03:00 MCC update. MCC is going to send 5 uA to the hall to calibrate the BCM. Setting left HRS to 1 GeV/c.

    04:00 MCC asks for BeO target to check ep-ion chamber. Set HRS down to 0.5 GeV/c to be sure we do not have a sharp elastic peak on a small part of the focal plane. MCC sends upto 8.6 uA -- beam spot is possibly visible there for someone with an (over?)active imagination.

    04:05 Request raster of 1 mm horizontal by 5 mm vertical. Spot++ does not work, possibly due to no trigger8. Bogdan suggests we try to lower pion rejector HV since signals are saturating.

    We could not quickly find a map of HV channel to pion rejector channel so we turn off card 9 channel 0 from 1505 to 0V. In replay we see pion rejector 2 channel left arm shower 17 is now off. Return card 9 channel 0 to 900 V (it was saturating). Turn off card 9 channel 1 to confirm. Pion rejector 2 channel left arm shower 18 is now off. Assume the next 10 channels on the card are sequential to shower 28.

    Turn off card 10 channel 0. Pion rejector 2 channel left arm shower 29 is now off. Turn off card 10 channel 5. Channel 16 of Pion Rejector 1 is missing.

    Turn off card 10 channel 6. Channel 17 is missing on Pion Rejector 1. Card 11 channel 0 is Pion Rejector 1 Channel 23 (as expected).

    05:00 Bogdan requests access to work on trigger. Turn off septum.

    07:00 out of access, set fields to 1 GeV/c, get beam back. Page Vince and Bob to come in later today. Start cycling quads for optics work. Simultaneously, go to empty target and ask for harp scan. Note sieve slit is out, so optics might start with multi-foil target.

    07:45 Set HRS to 1.166 per Vince's optics plan. For Q3 went to 1400 and turned on field regulation. For Q2, went to 1000 A, 0, 1000A, 0, 1000A, and turned on field regulation.